# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ 3.2 Estimation procedure
2:Obtain an initial estimate ${\bm{\theta}}^{0}$ by solving $\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})={\bf 0}$.
3:for $k$ in $1,\cdots\cdots,K$ do
4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q$.
5:     Update $\widehat{\bm{\theta}}^{k}$ by solving (11) with $\bm{\omega}^{k-1}$.
6:until convergence criterion $\|\widehat{\bm{\theta}}^{k}-\widehat{\bm{\theta}}^{k-1}\|<\delta$ satisfied.
7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}(\widehat{\bm{\omega}})$.
@@ 3.3 Examples: linear and logistic regression
Here, we give two examples to provide intuitions for POP-Inf estimation procedure. We start with estimating coefficients in a linear regression model where the resulting estimator has a closed form solution.
@@
We consider the ordinary least squares problem, where $\bm{\psi}(y,{\bf x};{\bm{\theta}})={\bf x}(y-{\bf x}^{\rm T}{\bm{\theta}})$. Therefore, the estimand is ${\bm{\theta}}_{0}=\mathbb{E}[{\bf X}{\bf X}^{\rm T}]^{-1}\mathbb{E}[{\bf X}Y]$. From equation (11), the POP-Inf estimator should be the solution to the equation
$\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}(y_{i}-{\bf x}_{i}^{\rm T}{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\omega}\odot{{\bf x}}_{i}\{\widehat{f}({\bf z}_{i})-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\omega}\odot{\bf x}_{i}\{\widehat{f}({\bf z}_{i})-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}={\bf 0},$ (20)
yielding a closed-form solution for POP-Inf linear regression
$\widehat{\theta}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
where $\widehat{\bm{\omega}}$ is a consistent estimate of optimal $\bm{\omega}_{0}$, which can be obtained by Algorithm 1.
The second example is logistic regression, which does not have a closed-form solution.
@@
For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$. Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
$\dfrac{1}{n}\sum_{i=1}^{n}\left\{-{\bf x}_{i}y_{i}+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})\right\}+\bm{\omega}\odot\left\{\dfrac{1}{N}\sum_{i=n+1}^{n+N}[-{{\bf x}}_{i}\widehat{f}({\bf z}_{i})+{{\bf x}}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]-\dfrac{1}{n}\sum_{i=1}^{n}[-{\bf x}_{i}\widehat{f}({\bf z}_{i})+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]\right\}={\bf 0}.$ (22)
This does not yield a closed form solution. To obtain an estimator, one can adopt Algorithm 1.
@@ 4 Relationship with Efficient Influence Function
To connect our method with the literature of semiparametric efficiency (Bickel et al.,, 1993; Tsiatis,, 2006), in this section, we state the efficient influence function (EIF) and the semiparametric efficiency bound for the estimation problem (1) (Proposition 1) and connect it with our proposed POP-Inf estimator.
To ease notation, we introduce a random variable $R$ to indicate the labeling; $r_{1}=\cdots=r_{n}=1$ and $r_{n+1}=\cdots=r_{n+N}=1$. In this section, suppose $\pi\equiv\hbox{pr}(R=1)=n/(n+N)$ is a fixed constant between zero and one. Note that the introduction of $R$ is for notational simplicity and we make no inference on it.
@@
Proof is contained in Appendix B.2.
Guided by Proposition 1, an EIF-based estimator would be the solution to the empirical equation,
@@
As stated in Corollary 2, this EIF-basd estimator is asymptotically normal and achieves the semiparametric efficiency bound.
@@
Under Conditions (C1) and (C3), assume $\mathbb{E}\{\sup_{{\bm{\theta}}\in{\bm{\Theta}}}\bm{\phi}(Y,{\bf X},{\bm{\theta}})\}<\infty$ and equation (24) has a unique solution, then as $n\to\infty$ and $N\to\infty$,
@@
If the nuisance function that appears in the EIF, in particular $\mathbb{E}(Y\mid{\bf x},{\bf z})$ here, can be correctly specified or well approximated, then an EIF-based estimator is feasible in practice, and it will achieve the semiparametric efficiency bound. However, this is nearly impossible in reality with limited knowledge of the data generation process or a small computational budget. Instead, it’s typical to construct an estimator with nuisance functions plugged in as their estimates. We refer to such estimators as “$*{}^{*}$-based”, because the influence function of the resulting estimator, denoted by $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$, may or may not be the efficient influence function.
In our problem, the nuisance function $\mathbb{E}(Y\mid{\bf x},{\bf z})$ can be estimated with $\widehat{f}({\bf z})$, which can be far from satisfactory even if the covariate of interest ${\bf x}$ is a subset of the predictor ${\bf z}$. Upon simple calculation, we can write the resulting $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$ as the solution to
$\dfrac{1}{n}\sum_{i=1}^{n}{\bf A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
Similar to the equation (11) that the POP-Inf estimator solves, the $*{}^{*}$-based estimator (26) solves a equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term utilizes predictions $\widehat{f}$ that is always mean zero. If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (26) corresponds to that equation weight equals to a vector of $1-\pi$. In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially the scalar one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$. The difference between (11) and (26) clarifies one of the main distinctions between the proposed POP-Inf estimator and an EIF guided estimator. The former seeks for variance reduction in a data-adaptive way, while the latter attains minimum variance at correct specifications. In real-world situations, the proposed POP-Inf estimator can be much more efficient in that it utilizes information in a data-adaptive fashion. Numeric comparisons are provided in Section 6.

### v4

@@ 3.1 PSPA estimator
2:Obtain the classical estimator $\widehat{\bm{\theta}}_{\text{C}}$ by solving $\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})={\bf 0}$.
3:Obtain the optimal weighting vector $\widehat{\bm{\omega}}^{\textnormal{opt}}=[\widehat{\omega}_{1}^{\textnormal{opt}},\dots,\widehat{\omega}_{q}^{\textnormal{opt}}]^{\rm T}$ by
$\widehat{\omega}_{j}^{\textnormal{opt}}=\min(\frac{[\widehat{\bf A}_{\text{C}}^{-1}\widehat{\bf M}_{4,{\text{C}}}\widehat{\bf A}_{\text{C}}^{-1}]_{jj}}{[\widehat{\bf A}_{\text{C}}^{-1}(\widehat{\bf M}_{2,{\text{C}}}+\rho\widehat{\bf M}_{3,{\text{C}}})\widehat{\bf A}_{\text{C}}^{-1}]_{jj}},1)\text{ for all }j\in 1,\dots,q,$
where $\widehat{\bf A}_{\text{C}}$, $\widehat{\bf M}_{1,{\text{C}}}$, $\widehat{\bf M}_{2,{\text{C}}}$, $\widehat{\bf M}_{3,{\text{C}}}$, $\widehat{\bf M}_{4,{\text{C}}}$ are sample analogs of ${\bf A}$, ${\bf M}_{1}$, ${\bf M}_{2}$, ${\bf M}_{3}$, ${\bf M}_{4}$ with $\widehat{\bm{\theta}}_{\text{C}}$ plugged in.
4:Obtain the PSPA estimator with a one-step update:
$\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}}}=\widehat{\bm{\theta}}_{\text{C}}-\left[\nabla\Psi_{\textnormal{{PSPA}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}(\widehat{\bm{\theta}}_{\text{C}})\right]^{-1}\Psi_{\textnormal{{PSPA}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}(\widehat{\bm{\theta}}_{\text{C}}),\text{ where }\nabla\Psi_{\textnormal{{PSPA}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}=\partial\Psi_{\textnormal{{PSPA}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}({\bm{\theta}})/\partial{\bm{\theta}}|_{{\bm{\theta}}=\widehat{\bm{\theta}}_{\text{C}}}$
5:Obtain the asymptotic variance of $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}}}$:
$\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}}^{\textnormal{opt}})=\widehat{\bf A}_{\textnormal{{PSPA}}}^{-1}[\widehat{\bf M}_{1,\textnormal{{PSPA}}}+{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})(\widehat{\bf M}_{2,\textnormal{{PSPA}}}+\rho\widehat{\bf M}_{3,\textnormal{{PSPA}}}){\rm diag}({\bm{\omega}})-2{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})\widehat{\bf M}_{4,\textnormal{{PSPA}}}]\widehat{\bf A}_{\textnormal{{PSPA}}}^{-1},$
where $\widehat{\bf A}_{\textnormal{{PSPA}}}$, $\widehat{\bf M}_{1,\textnormal{{PSPA}}}$, $\widehat{\bf M}_{2,\textnormal{{PSPA}}}$, $\widehat{\bf M}_{3,\textnormal{{PSPA}}}$, $\widehat{\bf M}_{4,\textnormal{{PSPA}}}$ are sample analogs of ${\bf A}$, ${\bf M}_{1}$, ${\bf M}_{2}$, ${\bf M}_{3}$, ${\bf M}_{4}$ with $\widehat{\bm{\theta}}_{\textnormal{{PSPA}}}$ plugged in.
6:PSPA estimator $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}}}$, standard error $\sqrt{\frac{\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}}$, $\alpha$-level confidence interval $\mathcal{C}_{\alpha,j}^{\textnormal{{{PSPA}}}}=(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}_{j}}\pm z_{1-\alpha/2}\sqrt{\frac{\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}})$, and (two-sided) p-value $2(1-\Phi(|\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPA}}_{j}}}{\sqrt{\frac{\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}})}|))$ for the $j$-th coordinate. Here, $\Phi$ is the CDF of the standard normal distribution.
3.2 Theoretical guarantees
Next, we establish the theoretical guarantees for our proposed estimator and algorithm. We modify the Theorem 1 to reflect that $\widehat{\bm{\omega}}$ in the algorithm is estimated from the data.
Corollary 1.
Suppose $\widehat{\bm{\omega}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\omega}}$ and conditions for Theorem 1 hold, then $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}$ and
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})-{\bm{\theta}})\stackrel{{\scriptstyle D}}{{\rightarrow}}N({\bf 0},{\bm{\Sigma}}({\bm{\omega}})).$ (16)
Proof of Corollary 1 is contained in Appendix B.2. In our algorithm, we substitute the sample analogs for ${\bf A}$, ${\bf M}_{1}$, ${\bf M}_{2}$, ${\bf M}_{3}$, ${\bf M}_{4}$ into $\omega_{j}^{\textnormal{opt}}$. Given that these sample analogs are typically consistent estimator, Slutsky’s theorem implies that $\widehat{\bm{\omega}}^{\textnormal{opt}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\omega}}^{\textnormal{opt}}$. Therefore, this additional condition is satisfied by our algorithm. The asymptotic normality in Corollary 1 guarantees the validity for inference for our algorithm by the following Corollary 2.
Corollary 2.
Suppose $\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\Sigma}}({\bm{\omega}})$ and conditions for Theorem 1 hold. Given confidence level $1-\alpha\in(0,1)$,
$\lim_{n\rightarrow\infty}\mathbb{P}(\theta_{j}\in\mathcal{C}_{\alpha,j}^{\widehat{\bm{\omega}}})\geq 1-\alpha,\ j=1\cdots,q,$ (17)
where $\theta_{j}$ is the $j$-th coordinate of the parameter and $\mathcal{C}_{\alpha,j}^{\textnormal{{{PSPA}}}}=(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}_{j}}(\widehat{\bm{\omega}})\pm z_{1-\alpha/2}\sqrt{\frac{\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}})_{jj}}{n}})$.
Proof of Corollary 2 follows from the asymptotic normality of the PSPA estimator in Corollary 1. Since we plug in the consistent estimator for each element in ${\bm{\Sigma}}({\bm{\omega}}^{\textnormal{opt}})$, Slutsky’s theorem implies that $\widehat{\bm{\Sigma}}(\widehat{\bm{\omega}}^{\textnormal{opt}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\Sigma}}({\bm{\omega}}^{\textnormal{opt}})$, and therefore the condition in Corollary 2 is also satisfied by our algorithm.
The above two corollaries ensure the validity of the inference for our algorithm. We next demonstrate that our method achieves element-wise variance reduction when compared to all baseline methods.
Proposition 1.
Suppose $\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{{PSPA}}}({\bm{\omega}})-{\bm{\theta}})\stackrel{{\scriptstyle D}}{{\rightarrow}}N({\bf 0},{\bm{\Sigma}}({\bm{\omega}}))$, given ${\bm{\omega}}^{\textnormal{opt}}$ defined by equation 14, ${\bm{\Sigma}}({\bm{\omega}}^{\textnormal{opt}})_{jj}\leq{\bm{\Sigma}}({\bm{\omega}})_{jj}$ for all $j\in\{1,\cdots,q\}.$
Proof of Proposition 1 is contained in Appendix B.3. By setting the weighting vector ${\bm{\omega}}$ to ${\bm{\omega}}^{\textnormal{C}}=[0,\dots,0]^{\rm T}$ or ${\bm{\omega}}^{\textnormal{{PPI}}}=[1,\dots,1]^{\rm T}$, our method reduces to classical method and PPI, respectively. Therefore, PSPA guarantees an element-wise reduction in the asymptotic variance compared to both the classical method and PPI, for arbitrary ML models.
@@ 3.3 Examples: linear and logistic regression
Here, we present two examples to illustrate the intuitions behind the PSPA estimation procedure. We begin with the estimation of coefficients in a linear regression model, where the resulting estimator is expressed in a closed form.
@@
We consider the ordinary least squares problem, where $\bm{\psi}(y,{\bf x};{\bm{\theta}})={\bf x}(y-{\bf x}^{\rm T}{\bm{\theta}})$. Therefore, the estimand is ${\bm{\theta}}_{0}=\mathbb{E}[{\bf X}{\bf X}^{\rm T}]^{-1}\mathbb{E}[{\bf X}Y]$. From equation (11), the PSPA estimator is the solution to the equation
$\small\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}(y_{i}-{\bf x}_{i}^{\rm T}{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\rm diag}({\bm{\omega}}){{\bf x}}_{i}\{\widehat{f}-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}-\dfrac{1}{n}\sum_{i=1}^{n}{\rm diag}({\bm{\omega}}){\bf x}_{i}\{\widehat{f}-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}={\bf 0},$ (18)
yielding a closed-form solution for PSPA linear regression
$\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}=$$\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}$ (19)
$\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}\right]\right\},$ (20)
where $\widehat{\bm{\omega}}^{\textnormal{opt}}$ is a consistent estimator of optimal ${\bm{\omega}}^{\textnormal{opt}}$, obtainable through Algorithm 1.
The second example is logistic regression, which lacks a closed-form solution but can be solved using Algorithm 1.
@@
For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$. Using estimation equation (11), the PSPA estimator $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}$ is the solution to the equation
$\small\dfrac{1}{n}\sum_{i=1}^{n}\left\{-{\bf x}_{i}y_{i}+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})\right\}+{\rm diag}({\bm{\omega}})\left\{\dfrac{1}{N}\sum_{i=n+1}^{n+N}[-{{\bf x}}_{i}\widehat{f}({\bf z}_{i})+{{\bf x}}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]-\dfrac{1}{n}\sum_{i=1}^{n}[-{\bf x}_{i}\widehat{f}+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]\right\}={\bf 0}$ (21)
@@ 4 Relationship with Efficient Influence Function
To connect our method with the literature on semiparametric efficiency (Bickel et al.,, 1993; Tsiatis,, 2006), in this section, we state the efficient influence function (EIF) and the semiparametric efficiency bound for the estimation problem (1) (Proposition 2) and connect it with our proposed PSPA estimator. To ease notation, we introduce a random variable $R$ to indicate the labeling; $r_{1}=\cdots=r_{n}=1$ and $r_{n+1}=\cdots=r_{n+N}=1$. In this section, suppose $\pi\equiv\hbox{pr}(R=1)=n/(n+N)$ is a fixed constant between zero and one. Note that the introduction of $R$ is for notational simplicity and we make no inference on it.
@@
The proof is contained in Appendix B.4. Guided by Proposition 2, an EIF-based estimator would be the solution to the empirical equation,
@@
With the derived EIF, the EIF-based estimator is asymptotically normal and achieves the semiparametric efficiency bound by section 8.4 of Molenberghs et al., (2014).
@@
Under Conditions (C1) and (C3), assume
$\mathbb{E}\{\sup_{{\bm{\theta}}\in{\bm{\Theta}}}\bm{\phi}(Y,{\bf X},{\bm{\theta}})\}<\infty$ and equation (23) has a unique solution, then as $n\to\infty$ and $N\to\infty$,
@@
If the nuisance function $\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}$ that appears in the EIF can be correctly specified or well approximated, then an EIF-based estimator will achieve the semiparametric efficiency bound. However, this is nearly impossible in reality with limited knowledge of the data generation process or a small computational budget. Instead, it’s typical to construct an estimator with nuisance functions plugged in as their estimates. We refer to such estimators as “EIF-based”, because the influence function of the resulting estimator, denoted by $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$, may or may not be the efficient influence function.
In our setting, the nuisance function $\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}$ can be estimated with $\mathbb{E}\{\bm{\psi}(\widehat{f},{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}$. Upon simple calculation, we can write the resulting $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$ as the solution to
$\small\dfrac{1}{n}\sum_{i=1}^{n}{\bf A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (25)
Similar to the equation (11) that the PSPA estimator solves, the EIF-based estimator (25) solves an equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term that utilizes predictions $\widehat{f}$, which always have a mean of zero. If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (25) corresponds to that equation weight equals to a vector of $1-\pi$. In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially the scalar one, the EIF-based estimator is the PSPA estimator with weight being $1-\pi=N/(n+N)$. The difference between (11) and (25) clarifies one of the main distinctions between the proposed PSPA estimator and an EIF-based estimator. The former seeks variance reduction in a data-adaptive way, while the latter attains minimum variance at correct specifications. In real-world situations, the proposed PSPA estimator can be much more efficient in that it utilizes information in a data-adaptive fashion. Numeric comparisons are provided in Section 6.